OpenStudy (anonymous):

what is dy/dx of x^3-xy+y^3=1 and how?

OpenStudy (anonymous):

3*x^2 - (y + xdy/dx) + 3y^2 * dy/dx = 0 3x^2 - y - xdy/dx + 3y^2 * dy/dx = 0 dy/dx * (3y^2 - x) = y - 3x^2 dy/dx = (y - 3x^2)/(3y^2 - x)

OpenStudy (anonymous):

Thanks a ton!